OpenStudy (anonymous):

Find the value of 2 times the square root of the quantity 3 over 25.. Options are below 1 and 1 half. 2 times the square root of 3, all over 5. 2 times the square root of 3, all over 25. 6 over 25.

OpenStudy (anonymous):

Please Help

OpenStudy (anonymous):

I think it's the third one.

OpenStudy (anonymous):

2\[\sqrt{3/25}\]

OpenStudy (anonymous):

Wait the last one was wrong

OpenStudy (anonymous):

Noemi how did u get that?

OpenStudy (anonymous):

okay then post the last one.

OpenStudy (anonymous):

no I meant what I posted

OpenStudy (anonymous):

I believe it's number 2

OpenStudy (anonymous):

2*sqrt(3)/5

OpenStudy (anonymous):

How did you get the trog? I really appreciate this thank you guys

OpenStudy (anonymous):

Here's my calculations first, simplify the square root (sqrt) sqrt(3/25) = sqrt(3)/sqrt(25) (property of exponents) sqrt(25)= 5 --> sqrt(3)/sqrt(25) = sqrt(3)/5 sqrt(3) is a radical, so we can't simplify it further, so just multiply the whole thing by 2

OpenStudy (anonymous):

Sorry, giving the final result 2*sqrt(3)/5

OpenStudy (anonymous):

True, I agree with Trog now!

OpenStudy (anonymous):

Okay thank you so much! I am about to post another one if you could help me with that too?